OpenStudy (anonymous):

The number of solutions that are possible for the inequality \[\left[ x-5 \right] +\left| x -1\right|<2\] is

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

did this one yesterday there is no solution

OpenStudy (anonymous):

where....

OpenStudy (anonymous):

The answer is 0 solution u r correct but how????

OpenStudy (anonymous):

there are three intervals to consider \(x<1\) and \(1<x<5\) and \(x>5\) if \(1<x<5\) then \(|x-1|=x-1\) and \(|x-5|=5-x\) so \(|x-5|+|x-1|=5-x+x-1=4\) in that interval and clearly 4>2

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

if \(x>5\) then \(|x-1|=x-1, |x-5|=x-5\) and so \(|x-5|+|x-1|=x-5+x-1=2x-6\) in that interval now if \(x>5\) then \(2x-6>4\) so it is larger than 4 in that interval as well, and if it is larger than 4 it is not less than 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the last interval to consider is if \(x<1\) in which case \(|x-1|=1-x\) and \(|x-5|=5-x\) and therefore .... you can finish and conclude that if \(x<1\) your expression is still larger than 4, so it is never less than two

OpenStudy (anonymous):

then for each case condradict

OpenStudy (anonymous):

yes, i left the last case up to you, but it is pretty much like the second one

OpenStudy (anonymous):

here is a nice visual confirmation that this expression is always \(\geq 4\) http://www.wolframalpha.com/input/?i=y%3D |x-5|%2B|x-1|

OpenStudy (anonymous):

you'll have to cut and past, absolute values reek havoc with links

OpenStudy (anonymous):

so .... how do solution become 0

OpenStudy (anonymous):

but the answer should be 0 dude...............

OpenStudy (anonymous):

@glgan1 help lzz

OpenStudy (anonymous):

@experimentX can u explain this more clearly....

OpenStudy (anonymous):

how it condradicts can u show that

OpenStudy (experimentx):

sorry ignore my last post ..

OpenStudy (anonymous):

ok

OpenStudy (experimentx):

it's more clearly explained here http://openstudy.com/study#/updates/4fbf6570e4b0dd370c772b87 the lowest value of LHS is 4 ... 4 cannot be greater than 2... which is a contradiction

OpenStudy (anonymous):

if i solutions condradicts....then others.....

OpenStudy (experimentx):

what exactly are you expecting to be the solution??

OpenStudy (anonymous):

i want to know why the solution become 0

OpenStudy (experimentx):

the solution is no zero ... simply there is not any solution

OpenStudy (anonymous):

@experimentX anyway a many thanzz to u .... becozz u r the only person who reponded to my message quickly and help me thanzzzz a billion!!

OpenStudy (anonymous):

sorry helped me!!

OpenStudy (experimentx):

np ...

OpenStudy (anonymous):

@heena

OpenStudy (anonymous):

@open_study1 wat happen ur qn is already answered i guess

OpenStudy (anonymous):

i did nt understand............